# NTSEBench: Cognitive Reasoning Benchmark for Vision Language Models

### v1

@@ 4.1 Results
Table 4: Multi-modality Results: Zero-shot and Few-shot performance of different VLM models across various Text+Vision categories. We report results using all three different modelling strategies proposed to handle multimodality data, i.e., Interleaved, Image Only and Standard VQA. italics font for proprity models, i.e., money or API access is required to run these models. Note: (* )In some models, a common issue arises when a model refrains from providing a response due to safety concerns, often stemming from misinterpretation of the image’s intent.e.g thinking it as CAPTCHA.
@@
Question category analysis. The results from Table 3 reveal that while LLMs generally perform better on the text-only subset of NTSEBench, the high standard deviation of 11 indicates significant variability in model performance across different question types. This variability may stem from overlaps between NTSEBench and other open-source datasets, suggesting that there are still areas where LLMs exhibit limitations in reasoning capabilities. This is especially evident in Alphabet Test(ALP) category and Mathematical Operations(MTO) category, where the accuracy is more than two standard deviation away from mean accuracy of the model. This could also be attributed to the potential difficulty of these question types but that analysis has been left for future work.
@@
VLMs have shown poor performance across all categories for multi-modal questions, with the best-performing model achieving correct answers only 34% of the time. Even the standard deviation for accuracy of the best model is 9, indicating that VLMs struggle more with certain question categories than others. Specifically, we observe that VLMs perform notably poorly on categories such as DOT (Dot Problems), NVS (Non-Verbal Series), and NVO (Non-Verbal Odd One Out). These categories require identifying correlations or patterns between multiple images, or recognizing emerging patterns in a sequence of images. This task is akin to identifying similar and evolving patterns in different parts of an image, and predicting the next possible pattern. While vision models excel at recognizing existing patterns, they struggle with predicting new patterns. Improving VLM model performance on these tasks may necessitate exploring different architectures and training methodologies in the future.
@@
NTSEBench presents a challenging task for SOTA LLMs and VLMs. Based on the findings in Tables 3 and 4, it is evident that the proposed dataset presents a challenging task for all state-of-the-art LLM and VLM models. None of the open-source models achieve accuracy exceeding 45% on text-only questions and 18% on multi-modal questions, with propriety models achieving 62% and 42% accuracy respectively. Many of the models tested didn’t every reach random selection baseline of 24.52%(261 question had 5 options and 2467 question has 4 options), indicating that current LLMs and VLMs have a long way to go to attain better accuracy on these cognitive reasoning questions. Moreover, these findings highlight the potential for future advancements in addressing textual and visual reasoning questions for a variety of categories.
@@ 5 Related works
Texual and Multimodal Reasoning dataset. There exist multiple datasets to test domain specific(math, science, medicine) QA and reasoning abilities of LLMs and VLMs knowledge such as SciBench Wang et al. (2023b), SciEvalSun et al. (2024), MMMU Yue et al. (2024), MathVista Lu et al. (2023), JEEbench Arora et al. (2023), MathVerse Zhang et al. (2024a), OlympiadBench He et al. (2024) and many others on based on real world images or other domains He et al. (2020); Soni et al. (2022); Liu et al. (2023); Thrush et al. (2022); Li et al. (2023a). Most existing visual and multi-modal reasoning datasets in the literature are domain-specific M3Exam Zhang et al. (2024b), RAVENZhang et al. (2019), or they involve reasoning about real-world images Liu et al. (2023); Wang et al. (2024); Thrush et al. (2022); Li et al. (2023a) with basic commonsense reasoning questions such as CLEVR Johnson et al. (2017). However, current research has not thoroughly explored the capabilities of large models in addressing cognitive reasoning problems for both textual and multi-modal data. Datasets such as Liu et al. (2023); Wang et al. (2024); Masry et al. (2022) may explore mathematical reasoning in visual context but don’t allow for multiple images within the same prompt or reasoning across images, where finding a pattern within a sequence of images is required which is the case in many categories of NTSEBench like Paper Folding and Cutting, Embedded Figure, Figure Partition, Mirror/Water images etc. NTSEBench is different from all already existing datasets in the literature because it explicitly focuses on testing cognitive reasoning abilities of large deep learning models.
@@
Zero shot and Few shot prompting engineering for textual and multi-modal input. Our work is also related to prompting design and prompt engineering for LLMs Brown et al. (2020); Chen et al. (2023); Gupta et al. (2023); Khot et al. (2022); Wei et al. (2022); Sahoo et al. (2024); Ali et al. (2024) and VLMs gu2023survey; Xu et al. (2024); Bai et al. (2023); Liu et al. (2024); Dai et al. (2024).There are numerous studies on multi-modal and vision Chain-of-Thought (CoT) prompting Zhang et al. (2023); Shao et al. (2024). However, these studies do not explore the potential of state-of-the-art vision-language models (VLMs) to process multiple images simultaneously. This limitation was often due to the smaller context windows of older models. However, newer models now support this functionality and can even handle few-shot examples for CoT prompting. We explore three distinct prompting strategies made possible by the extended context windows of large vision-language models (VLMs).
@@ 6 Conclusion and Future Work
We have developed a new dataset, NTSEBench, specifically designed to assess the advanced analytical and logical reasoning capabilities of large deep learning models (LLMs and VLMs). We also propose four distinct modeling strategies for handling multi-modal data (text and images) across different question types in NTSEBench. These strategies enable us to conduct a fair and comprehensive comparison between closed and open-source models using both zero-shot and few-shot scenarios. Our findings indicate that both LLMs and VLMs struggle with advanced visual reasoning tasks, with VLMs performing particularly poorly on multi-modal questions compared to LLMs on textual questions. Additionally, proprietary models consistently outperform open-source models, correctly predicting twice as many questions. Overall, our results underscore that NTSEBench poses significantly greater challenges for state-of-the-art LLMs and VLMs.
Future directions. (a) We did not used image-to-text conversion because most of images in NTSEBench are not real-world images, and caption generation models are typically trained on real-world images, leading to a domain shift. This results in subpar captions that are neither relevant nor self-explanatory. Because LLMs performed significantly better at reasoning task, we intend to utilize caption generators to investigate whether text-only models, incorporating image descriptions generated by caption models, can outperform VLMs. (b)Given the limited data available for cognitive reasoning questions, we plan to use data augmentation strategies to increase the number of samples. (c) We also aim to investigate the impact of shuffling multiple-choice question (MCQ) options on performance in reasoning tasks.
@@
While our dataset is new and sourced for different sources when compared to dataset already present in the literature, there still might be some overlap in reasoning questions especially in textual reasoning. So all the dataset instances might not be independent and exclusive. This dataset is solely created in English, so no other languages are represented, therefore we cannot analyze whether language variations can have a significant impact on the reasoning capabilities of these large models. Our modelling strategies were limited to zero shot and few shot COT prompting. We did not evaluate whether finetuning these large models on few examples from each of the categories would further improve results. This was due to limitation of both GPU resources and large cost of fine-tuning for Proprietary models. We plan to address all these limitation in future extension of this work.
@@ Acknowledgements
Research was sponsored by the Army Research Office and was accomplished under Grant Number W911NF-20-1-0080. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. This work was partially funded by ONR Contract N00014-19-1-2620. We extend our gratitude to the annotators who verified our flowcharts and corresponding question answer pairs. Lastly, we extend our appreciation to the reviewing team for their insightful comments.
Appendix A Example Appendix
A.1 Problem Categories
Problem categories with description are shown in Table 5.

### v2

@@ 4 Results and Analysis
Table 3: Multi-modality Question Results: Zero-shot and Few-shot performance of different VLM models across various Text+Vision categories. We report results using all three different modelling strategies proposed to handle multimodality data, i.e., Interleaved, Image Only and Standard VQA. italics font for proprity models, i.e., money or API access is required to run these models. Note: (* )In some models, a common issue arises when a model refrains from providing a response due to safety concerns, often stemming from misinterpretation of the image’s intent.e.g thinking it as CAPTCHA.
Results for multi-modality question are reported in Table 3. Results are reported using three different modelling strategies listed in the section above namely Interleaved, Image Only and Standard VQA. Columns in the table refer to different categories of visual text reasoning questions in NTSEBench such as DIR (direction sense), VEN (venn diagrams), TIM (time and clock), MIS (missing character), NVS (non-verbal series), NVO (non-verbal odd one out), NVA (non-verbal analogy), INC (incomplete figure), MIR (mirror, water and images), CUB (cube and dice), PAP (paper folding & cutting), EMB (embedded figure), FIG (figure partition), DOT (dot problem). Option bias ablation results for Gemini-1.5-pro and GPT-4o models are shown in Appendix Tables 8,  9,  10 and  11.
Proprietary models outperform Open Source models. From results in table 2 and 3, we can observe that proprietary models, such as Gemini Pro 1.5, GPT-4o and o1-preview, outperform other open source models in nearly every category question category. Especially o1-preview, which operates by utilizing internal chain-of-thought reasoning, allowing it to analyze complex problems step-by-step before arriving at a solution. Proprietary models demonstrate nearly double the accuracy of open-source models on NTSEBench questions, both in text-based and multi-modal tasks, across all modeling strategies outlined in the previous section. Notably, Gemini Pro 1.5 consistently outperforms GPT-4o across most strategies, excelling in both text and multi-modal question types. For open-source models, LLaVA-OneVision-72b-ov-chat excels in text and multi-modality tasks, achieving SOTA results in some vision categories. Ovis1.6-Gemma2-9B, while smaller, outperforms GPT-4o and LLaVA-OneVision in several vision categories, though its overall performance is still lower.
Modelling Strategy is important. For text-only questions, the Standard QA strategy clearly outperforms Image Only modeling. Introducing burden of doing OCR on top of reasoning tends to confuse models or exacerbate the difficulty of the task. This effect is particularly noticeable with smaller open-source models, which struggle to accurately extract characters and integrate them into context. However, proprietary models like GPT-4o and Gemini-Pro still achieve superior results using Image Only compared to open-source models employing Standard QA or text-only processing alone. Few-shot results present a mixed picture: while models like Mixtral and GPT-4o show improved performance with added exemplars, others experience a significant decline. Tables 2 and 3 show that proprietary models generally experience a smaller performance drop than open-source models in such scenarios.
Interleaving text and images performs better than Standard VQA and Image Only strategy for most categories. This underscores the importance of presenting text and images separately and in a more detailed manner, providing appropriate context, and treating the image and text as distinct entities. Our results in Table 3 show that this approach significantly improves Vision-and-Language Models (VLMs), as demonstrated by the superior performance of open-source models like LLaVA-OneVision-72b and Ovis-9b, which handle interleaved text and images more effectively than others. Table 3 shows that few-shot prompting consistently underperforms compared to zero-shot for all VLM models, suggesting that adding exemplars may confuse large models, hindering their focus on logical reasoning.
Multi-modal reasoning is significantly harder. Comparing the best and worst performing models in Tables 2 and 3, it’s evident that multi-modal reasoning is considerably more challenging than textual reasoning for current state-of-the-art VLMs. Multi-modal questions see less than 45% accuracy, while the best model for textual reasoning, o1-preview, exceeds 80%. Even fast models like Gemini 1.5 pro achieve over 60% accuracy in textual reasoning, highlighting the gap between VLMs and LLMs and the need for better architectures and datasets for VLMs.
@@
Question category analysis. The results from Table 2 reveal that while LLMs generally perform better on the text-only subset of NTSEBench, the high standard deviation (11.04) indicates significant variability in model performance across different question types. This variability may stem from some overlap between NTSEBench and other open-source datasets, suggesting that there are still areas where LLMs exhibit limitations in reasoning capabilities. This is especially evident in Alphabet Test(ALP) category and Mathematical Operations(MTO) category, where the accuracy is more than two standard deviations away from the mean accuracy of the model. This could also be attributed to the potential difficulty of these question types however that analysis has been left for future work.
@@
VLMs have shown poor performance across all categories for multi-modal questions, with the best-performing model achieving correct answers only 42% of the time. Even the standard deviation for accuracy of the best model is 9, indicating that VLMs struggle more with certain question categories than others. Specifically, we observe that VLMs perform notably poorly on categories such as DOT (Dot Problems), NVS (Non-Verbal Series), and NVO (Non-Verbal Odd One Out). These categories require identifying correlations or patterns between multiple images or recognizing emerging patterns in a sequence of images. This task is akin to identifying similar and evolving patterns in different parts of an image and predicting the next possible pattern. While vision models excel at recognizing existing patterns, they struggle with predicting new patterns.
@@
NTSEBench presents a challenging task for SOTA LLMs and VLMs. Based on the findings in Tables 2 and 3, it is evident that the proposed dataset presents a challenging task for all state-of-the-art LLM and VLM models. None of the open-source models achieve accuracy exceeding 50% on text-only questions and 35% on multi-modal questions, with propriety models achieving 62% and 42% accuracy respectively. Although o1-preview has created its own category and got more than 80% accuracy on text only data, it can not do multi-modal reasoning. Many of the models tested didn’t even reach random selection/guess baseline of 24.52%(261 question had 5 options and 2467 question has 4 options), indicating that current LLMs and VLMs struggle with cognitive reasoning questions.
Human vs Model. Preliminary human evaluation results shown in appendix section A.4.1 show that average human accuracy for textual and multi-modal questions is more than 80%, much greater than 62% (text) and 42% (visual) for best performing propriety model. Although the o1-preview model have achieved near human accuracy on textual reasoning, it is significantly slower than other models. These findings suggest potential for future advances in tackling diverse textual and visual reasoning questions.
Options Bias Ablation. From Gemini 1.5 Pro Results in Table  8 and  9 we can see that aggregate performance for both multi-modal and text question is affected by the position of the correct option(variation in performance ranging from -4 to +6 percent when compared to random for text and -5 to +5 percent for multi-modal categories). We can make similar observation on GPT-4o (tables  10 and  11), however, GPT-4o has smaller variation in performance than Gemini Pro. These variations suggest that the model’s responses to logical questions may be either memorized or guessed, indicating a potential limitation in LLMs and VLMs. Even o1-preview, as shown in Table 12, exhibits performance variations akin to those of GPT-4o, suggesting that advanced reasoning models are similarly affected by bias.
Error Analysis. We manually conducted an error analysis of 260 questions (10 from each question category) for Gemini 1.5 pro, and we identified distinct patterns in reasoning and error categorization. We have categorised errors based on the cognitive dimensions outlined in section 2. The Sankey diagram in figure 2 illustrates how errors across various question categories correspond to specific error types. A key observation is that many errors arise from Pattern Recognition failures, especially in categories like Alphabet Tests, Non-Verbal Analogy, and Series questions, where the model struggled with recurring patterns and sequence shifts, highlighting challenges in complex pattern-based reasoning. We also noted frequent errors in Spatial Reasoning and Logical Deduction tasks, particularly in spatial or diagrammatic questions like Cube and Dice, Embedded Figure, and Paper Folding & Cutting. These questions often require pattern recognition, shape manipulation, or deducing logical relations from limited visual data. The figure shows that errors in Quantitative Analysis were common in numerical questions like Time and Clock and Mathematical Operations, indicating the model excels in simpler tasks but struggles with complex number sequences and operations. The error distribution reveals key insights into the model’s strengths and weaknesses, guiding future improvements. A detailed error analysis for each category is included in Appendix  A.3.
Figure 2: Error Analysis. Overview of errors Gemini 1.5 Pro makes across different question categories.
@@ 5 Related works
Texual and Multimodal Reasoning dataset. There exist multiple datasets to test domain specific textual (math, science, medicine) QA and reasoning abilities of LLMs and VLMs knowledge such as SciBench Wang et al. (2023b), SciEvalSun et al. (2024), MMMU Yue et al. (2024), MathVista Lu et al. (2023), JEEbench Arora et al. (2023), MathVerse Zhang et al. (2024a), OlympiadBench He et al. (2024) and many others on based on real world images or other domains He et al. (2020); Soni et al. (2022); Liu et al. (2023); Thrush et al. (2022); Li et al. (2023). Even most existing visual and multi-modal reasoning datasets in the literature are domain-specific M3Exam Zhang et al. (2024b), RAVEN Zhang et al. (2019), or they involve reasoning about real-world images Liu et al. (2023); Wang et al. (2024); Thrush et al. (2022); Li et al. (2023) with basic commonsense reasoning questions such as CLEVR Johnson et al. (2017). However, current research has not thoroughly explored the capabilities of large models in addressing cognitive/critical reasoning problems for both textual and multi-modal data. NTSEBench is different from all already existing datasets in the literature because it explicitly focuses on testing cognitive reasoning abilities of large deep learning models. While benchmarks like BBH Suzgun et al. (2022) focus on text-based logical deduction and direction sense, NTSEBench offers a more holistic assessment by integrating both textual and visual elements. For example, it includes categories that combine text with visual answers, such as Venn diagrams and figure partition questions, which BBH does not cover. NTSEBench includes areas overlooked by other benchmarks, such as series, coding-decoding, and blood relations. With 2,728 multiple-choice questions, it provides a broader evaluation than the 770 puzzles in BBH.
Furthermore, benchmarks like Jiang et al. (2024b) emphasize visual abstraction and reasoning but lack the integration of text and visual elements found in NTSEBench. Our dataset also goes beyond Raven’s Progressive Matrices (RPMs) Zhang et al. (2019) and related visual analogy problems, which are central to Małkiński and Mańdziuk (2023), by incorporating a wider array of cognitive tasks. This includes diverse categories such as coding-decoding, number and ranking, blood relations, and mathematical operations. NTSEBench provides a fine-grained analysis of model performance across 26 distinct problem categories and investigates various modeling strategies, setting it apart from other benchmarks. Datasets like Wang et al. (2024); Liu et al. (2023); Masry et al. (2022) explore mathematical reasoning in visual contexts but do not permit multiple images in the same prompt or reasoning across images. This is crucial for identifying patterns in categories of NTSEBench, such as Paper Folding and Cutting, Embedded Figure, Figure Partition, and Mirror/Water images.
While ARC Chollet (2019) also targets general human cognitive understanding, its questions are primarily focused on visual tasks such as pattern completion, interpolation, denoising, and simple object counting. These tasks are relatively straightforward and often lack the need for complex reasoning. In contrast, NTSEBench spans 26 cognitive skills and integrates both textual and visual elements. By leveraging NTSE exam questions, it offers real-world relevance and facilitates quantitative evaluation through a multiple-choice format. Additionally, NTSEBench enables fine-grained performance analysis across cognitive dimensions, emphasizing knowledge integration and reasoning, which more closely mirrors human cognitive processes. These attributes make NTSEBench a more comprehensive and effective tool for evaluating AI’s cognitive capabilities.
Language Writ Large Harnad (2024) delves into the theoretical foundations of large language models (LLMs) like ChatGPT, offering insights into their unexpected capabilities through the analysis of technology dialogues. In contrast, NTSEBench provides quantitative performance metrics for LLMs across a wide range of reasoning tasks. While LogicVista focuses on logical reasoning in visual contexts through five distinct tasks, NTSEBench covers a broader spectrum, spanning 26 categories that include both visual and textual reasoning. MMMU, on the other hand, spans six disciplines and 30 college subjects, providing a broad overview of subject-specific knowledge. In contrast, NTSEBench offers a deeper, more focused evaluation of cognitive reasoning skills, making it an essential tool for assessing the reasoning capabilities of LLMs.
@@
Zero shot and Few shot prompting engineering for textual and multi-modal input. Our work is also related to prompting design and prompt engineering for LLMs Brown et al. (2020); Chen et al. (2023); Gupta et al. (2023); Khot et al. (2022); Wei et al. (2022); Sahoo et al. (2024); Ali et al. (2024) and VLMs Xu et al. (2024); Bai et al. (2023); Liu et al. (2024); Dai et al. (2024).There are numerous studies on multi-modal and vision Chain-of-Thought (CoT) prompting Zhang et al. (2023); Shao et al. (2024). These studies overlook the ability of state-of-the-art vision-language models (VLMs) to process multiple images simultaneously, a limitation of older models with smaller context windows. Newer models can handle few-shot examples for CoT prompting, and we explore three prompting strategies enabled by their extended context windows.
@@ 6 Conclusion and Future Work
We developed a new dataset, NTSEBench, to assess the advanced analytical and logical reasoning capabilities of large deep learning models (LLMs and VLMs). We also propose four distinct modeling strategies for handling multi-modal data (text and images) across different question types in NTSEBench. These strategies enable us to conduct a fair and comprehensive comparison between closed and open-source models using both zero-shot and few-shot scenarios. Our findings show that both LLMs and VLMs struggle with advanced visual reasoning tasks, with VLMs performing worse on multi-modal questions than LLMs on textual ones. Proprietary models also consistently outperform open-source models, correctly predicting twice as many questions. Overall, our results underscore that NTSEBench poses significantly greater challenges for state-of-the-art LLMs and VLMs. Future directions. (a) Our results indicated that VLM models have difficulty predicting novel patterns, implying that addressing this challenge may involve either architectural modifications or the integration of generative models alongside VLMs for these question types.
(b) Given the limited data available for cognitive reasoning questions, we plan to use data augmentation strategies to increase the number of samples.
@@
While our dataset is new and sourced for different sources when compared to dataset already present in the literature, there still might be some overlap in reasoning questions, especially in textual reasoning. So all the dataset instances might not be independent and exclusive. This dataset is solely created in English, so no other languages are represented, therefore we cannot analyze whether language variations can have a significant impact on the reasoning capabilities of these large models. Our modelling strategies were limited to zero shot and few shot COT prompting. We did not evaluate whether fine-tuning these large models on a few examples from each of the categories would further improve results. This was due to limitation of both GPU resources and large cost of fine-tuning for Proprietary models. Finally, our human evaluation study involved only three annotators with undergraduate degrees, which may limit the generalizability of the results to the broader human population. We plan to address all these limitations in future extension of this work.
@@ Acknowledgements
Research was sponsored by the Army Research Office and was accomplished under Grant Number W911NF-20-1-0080. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. This work was partially funded by ONR Contract N00014-23-1-2364. We extend our gratitude to the annotators who verified our flowcharts and corresponding question answer pairs.
Appendix A Appendix
A.1 The Global Relevance of the NTSE Exam for AI
The NTSE exam, despite being conducted in India, holds significant relevance for the global AI community due to its unique focus on cognitive reasoning abilities rather than domain-specific knowledge. The NTSE’s diverse question categories as described in table  4, which assess various cognitive dimensions, offer a robust framework for evaluating AI models’ capacity to process information, recognize patterns, and solve problems across different domains. This emphasis on cognitive skills aligns with the pursuit of Artificial General Intelligence (AGI) , making the NTSE exam’s insights and challenges applicable to AI research and development on a global scale.
A.2 Additional Dataset Details
Table 1, reveals a skewed distribution across various question categories. Notably, textual categories like the Alphabet Test (ALP) and Mathematical Operations (MTO) contain 94 and 99 examples, respectively. In contrast, many vision-based categories are more challenging and typically include between 80 and 100 examples. For instance, the Non-Verbal Analogy (NVA) category, one of the most difficult, comprises 100 examples. While this skewed distribution could impact model performance, exploring its effects is beyond the scope of this manuscript and is left for future work.
Problem Sub categories. The above categories are further subdivided based on the different modality of input type(text or text-vision). Table 1 lists each sub-category for Text Only questions and Text+Vision questions, along with the respective count for each category. As in Table 1 most categories are represented well in the dataset.
Problem Categories
Problem categories in the dataset with description are shown in Table 4.